# Stereo Any Video: Temporally Consistent Stereo Matching

### v2

@@ Abstract
This paper introduces Stereo Any Video, a powerful framework for video stereo matching. It can estimate spatially accurate and temporally consistent disparities without relying on auxiliary information such as camera poses or optical flow. The strong capability is driven by rich priors from monocular video depth models, which are integrated with convolutional features to produce stable representations. To further enhance performance, key architectural innovations are introduced: all-to-all-pairs correlation, which constructs smooth and robust matching cost volumes, and temporal convex upsampling, which improves temporal coherence. These components collectively ensure robustness, accuracy, and temporal consistency, setting a new standard in video stereo matching. Extensive experiments demonstrate that our method achieves state-of-the-art performance across multiple datasets both qualitatively and quantitatively in zero-shot settings, as well as strong generalization to real-world indoor and outdoor scenarios.
@@ 1 Introduction
Video stereo matching is a task to reconstruct a 3D scene from a stereo camera. It estimates disparity by identifying the horizontal correspondences between rectified left and right images at each time step. Consistently and accurately capturing disparities is crucial for reconstruction and understanding of real-world 3D scenes, benefiting various downstream tasks such as autonomous driving, robotic navigation, and virtual/augmented reality.
@@
To address these, we propose Stereo Any Video, a robust framework for video-based stereo matching, without relying on camera pose or optical flow. Inspired by recent findings in video generation [35, 50], which emphasize the role of robust feature representations in ensuring temporal coherence, we take a new perspective to design the framework for consistency: feature robustness and stability. Specifically, during feature extraction, we integrate convolutional cues with a frozen monocular video depth foundation model [8] trained on million-scale data. This ensures robust feature maps with rich priors. For correspondence matching, we introduce an all-to-all-pair correlation, which enforces smooth and reliable matching by jointly considering bidirectional correspondences from both target and reference views. For cost aggregation, we propose temporal convex upsampling within a fully temporal-based recurrent unit, playing a key role in maintaining temporal consistency. As shown in Figure 1, our method demonstrates a strong ability to produce accurate and consistent disparities, benefiting the reconstruction for dynamic scenes.
The contributions of our work can be summarized as follows: (1) Stereo Any Video, a framework that integrates monocular video depth priors to achieve accurate and temporally consistent video stereo matching. (2) A novel all-to-all-pair correlation module and a temporal convex upsampling mechanism, ensuring robust and stable cost volume representations and aggregation. (3) State-of-the-art performance on diverse benchmarks with at least 15% improvements, setting a new standard for video stereo matching.
Figure 2: Overview of the proposed Stereo Any Video. Given a stereo video sequence as input, the model first extracts features and context information using trainable convolutional encoders and a frozen monocular video depth encoder (Video Depth Anything [8]). At each iteration, an all-to-all-pair correlation computes feature correlations, followed by an MLP encoder for compact representation. Disparities are iteratively refined using a 3D Gated Recurrent Unit (GRU), integrated with cost volume features, and upsampled via a temporal convex upsampling layer. This process is repeated within a cascaded pipeline to progressively recover full-resolution disparities.
@@ 2 Related Work
Video Stereo Matching. Video stereo matching especially in dynamic scenes presents unique challenges. Early work, such as [78], employed LSTMs with unsupervised learning for stereo videos but did not explicitly incorporate temporal information. CODD [31] introduced a per-frame stereo network, a motion network, and a fusion network designed for dynamic scenes. Neighboring-frame disparities are aligned and fused using estimated camera poses. TemporalStereo [76] extended this approach with a coarse-to-fine network that leverages past context to enhance predictions in challenging scenarios. Cheng et al. [11] developed a framework tailored for XR systems, reducing computational costs through temporal cost aggregation. TC-Stereo [73] further introduced a temporal disparity completion mechanism and a temporal state fusion module to exploit temporal information. These methods rely on auxiliary camera motion or pre-defined scene geometry to incorporate temporal information. DynamicStereo [26] introduced a transformer-based architecture that integrates temporal, stereo, and spatial attention. BiDAStereo [24] proposed an optical-flow-based alignment mechanism to propagate temporal information across the entire sequence. Building on this alignment, BiDAStabilizer [25] introduced a lightweight stabilizer network to improve the temporal consistency of existing image-based stereo matching methods.
@@ 3 Method
In this section, we introduce Stereo Any Video, a feed-forward video-based framework. Given a rectified stereo sequence $\{\mathbf{I}^{t}_{L},\mathbf{I}^{t}_{R}\}_{t\in(1,T)}\in\mathbb{R}^{H\times W\times 3}$, our goal is to estimate a temporally consistent sequence of disparity maps $\{\mathbf{d}^{t}\}_{t\in(1,T)}\in\mathbb{R}^{H\times W}$ aligned with the left view, where $T$ denotes the total number of frames. As shown in Figure 2, our method consists of three key components: feature extraction (Section 3.1), correlation (Section 3.2), and temporal cost aggregation (Section 3.3).
@@ 3.1 Feature Extraction with Foundation Priors
Unlike image-based tasks, maintaining temporal consistency is the primary challenge in video stereo matching. Recent studies in video generation [35, 50] indicate that substantial variations across different frames lead to unstable feature representations, negatively impacting temporal coherence. Therefore, ensuring feature stability across adjacent frames is crucial for achieving temporally consistent disparity estimation. Towards this, we leverage Video Depth Anything (VDA) [8] for feature extraction, which offers two key advantages: (1) trained on internet-scale data, VDA provides rich priors that can enhance disparity estimation in regions with ambiguous stereo correspondence; (2) unlike image-based monocular depth models [71, 72], VDA generates temporally stable features through an explicit temporal alignment process before the final prediction head, thereby preserving coherence across frames.
Besides, we investigate whether the depth maps from monocular video depth models can further improve stereo performance. We compare DepthCrafter [21], a representative monocular video model, with RAFTStereo [33], an image-based stereo model. Using open-source weights, we evaluate their performance on Sintel [3], as shown in Table 1 and Figure 3. Specifically, we first obtain relative depth maps from DepthCrafter (without scale and shift) and compute disparities using RAFTStereo on the same sequence. The relative depth maps are then aligned to RAFTStereo’s disparities, and objective metrics are computed against the ground truth. Additionally, we conduct a subjective user study (see supplementary material for details). As observed, while DepthCrafter produces visually plausible results—reflected in better human scores—its spatial and temporal errors are significantly larger than those of RAFTStereo. This suggests that its relative depth is “consistent but wrong”, making it unreliable for directly improving consistency within a stereo framework. Instead, we utilize its features as valuable priors.
@@
Table 1: Comparison of monocular video depth method DepthCrafter [21] and image-based stereo matching method RAFTStereo [33] on Sintel dataset [3].
@@
Figure 3: The qualitative comparison of DepthCrafter [21] and RAFTStereo [33] on Sintel dataset [3].
@@ 3.2 All-to-all-pair Correlation
At the $n$-th iteration for each time step, the right feature $\mathbf{F}_{R}$ is first warped toward the left feature $\mathbf{F}_{L}$ using the previous disparity ${d}_{n-1}$, producing $\mathbf{\hat{F}}_{R}$. Previous methods [33, 64] typically construct cost volumes $\mathbf{C}$ via all-pair correlation, employ a lookup mechanism, and search for potential matches within a local radius. In [29], this approach was extended to an alternating local 1D and 2D search:
$\mathbf{C}(x,y)=\Big{\langle}\mathbf{F}_{L}(x,y),\hat{\mathbf{F}}_{R}(x+r_{x},y+r_{y})\Big{\rangle},$ (1)
@@
$\mathbf{C}(x,y)=\Big{\langle}\mathbf{F}_{L}(x+r_{x},y+r_{y}),\hat{\mathbf{F}}_{R}(x+r_{x},y+r_{y})\Big{\rangle}.$ (2)
@@
Table 2: Zero-shot generalization results. All methods were trained only on SceneFlow [36]. The representative image-based and video-based stereo matching models are compared on Sintel [3], Dynamic Replica [26], Infinigen SV [25], and Virtual KITTI2 [4]. The weights and parameters are fixed for evaluation. Best and second best results are highlighted.
@@ 3.3 Temporal Cost Aggregation
At the $n$-th iteration, the all-to-all-pair correlation expands the potential matching points in the cost volumes $\mathbf{C}_{n}$, as in Equation 2, increasing the dimensionality from $\mathbb{R}^{H^{\prime}\times W^{\prime}\times(2r_{x}+1)(2r_{y}+1)}$ to $\mathbb{R}^{H^{\prime}\times W^{\prime}\times(2r_{x}+1)^{2}(2r_{y}+1)^{2}}$, where $H^{\prime}$ and $W^{\prime}$ is the height and width of feature. We first encode $\mathbf{C}_{n}$ into a compact representation $\mathbf{E}_{n}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times L}$ using a Multilayer Perceptron (MLP):
@@
While prior methods emphasize temporal consistency in correlation and cost aggregation, the upsampling process remains overlooked, despite its crucial role in achieving accurate and temporally consistent disparities. Towards this, we develop the commonly used convex upsampling approach [53] to a temporal variant. In Figure 4, each high-resolution disparity pixel (yellow box) is computed as a convex combination of its $27$ low resolution neighbors across the $3$ frames (green, blue, and faint yellow boxes). The network predicts the corresponding $3\times 3\times 3$ weights using 3D convolutions, enabling a temporally consistent upsampling:
@@ 3.4 Loss Function
Recent works have introduced OPW loss [60] and TGM loss [8] in monocular video depth estimation to enhance temporal consistency. However, our investigation finds that these losses impose a trade-off between accuracy and consistency without yielding substantial improvement in video stereo task. Thus, we opt for a purely image-based $l_{1}$ distance loss while still achieving strong temporal consistency, as follows:
@@ 4 Experiments
Table 3: Quantitative comparisons on Sintel [3] and KITTI Depth [57] where methods are allowed to train on any existing datasets. Model weights and parameters remain fixed during evaluation. The 10 training datasets include SceneFlow [36], Sintel [3], Falling Things [54], InStereo2K [1], Carla [14], AirSim [47], CREStereo dataset [29], Middlebury [45], ETH3D [46], and KITTI 2012/2015 [15]. Training datasets sharing the same domain as the test sets are highlighted. Abbreviations: SF – SceneFlow [36], DR – Dynamic Replica [26], ISV – Infinigen SV [25], M – Middlebury [45], K – KITTI 2012/2015 [15], KD – KITTI Depth [57], VK – Virtual KITTI2 [4].
    Method # Training Data Sintel Clean Sintel Final KITTI Depth
@@
RAFTStereo [33] 10 mixed (incl. Sintel, K) 0.95 9.63 1.18 1.68 13.41 2.44 0.41 7.96 0.51
IGEVStereo [64] SF + M 0.91 8.42 1.25 1.53 13.49 2.23 0.46 9.12 0.60
Selective-IGEV [58] SF + M 0.88 9.04 1.16 1.64 13.78 2.14 0.48 9.66 0.60
@@
DynamicStereo [26] SF + DR + ISV + KD 0.72 8.19 1.36 1.38 11.84 3.38 0.42 8.21 0.57
BiDAStereo [24] SF + DR + ISV + KD 0.75 8.00 1.54 1.33 11.34 1.97 0.42 8.33 0.58
Stereo Any Video (ours) SF + DR + ISV + VK 0.55 6.44 0.89 0.99 9.09 1.48 0.35 6.05 0.61
@@ 4.1 Benchmark Datasets and Metric
Dataset. Our work focuses on videos captured with moving cameras, making commonly used image benchmarks unsuitable, such as Middlebury [45], ETH3D [46], and KITTI 2012/2015 [15] (two frames per scene). To evaluate our method, we use four synthetic and two real-world stereo video datasets, all featuring dynamic scenes. Sintel [3] is a synthetic dataset comprising movie sequences available in both clean and final passes. Dynamic Replica [26] is a synthetic indoor dataset featuring non-rigid objects, such as animals and people. Infinigen SV [25] is a synthetic dataset focusing on outdoor natural environments. Virtual KITTI2 [4] is a synthetic dataset simulating outdoor driving scenarios. KITTI Depth [57] is a real-world outdoor dataset captured for autonomous driving applications, with sparse depth maps obtained from a LiDAR sensor. South Kensington SV [25] is a real-world dataset capturing daily scenarios from a stereo camera for qualitative evaluation.
@@ 4.3 Evaluation
Table 4: Ablation study on the foundation priors, correlation, cost volume encoder, convex upsampling, attention, and inference settings of our model on Sintel [3] and Dynamic Replica [26]. For all metrics, lower values indicate better performance. The approach used in our final model is underlined.

### v3

@@ Abstract
This paper introduces Stereo Any Video, a powerful framework for video stereo matching. It can estimate spatially accurate and temporally consistent disparities without relying on auxiliary information such as camera poses or optical flow. The strong capability is driven by rich priors from monocular video depth models, which are integrated with convolutional features to produce stable representations. To further enhance performance, key architectural innovations are introduced: all-to-all-pairs correlation, which constructs smooth and robust matching cost volumes, and temporal convex upsampling, which improves temporal coherence. These components collectively enhance robustness, accuracy, and temporal consistency, establishing a new standard in video stereo matching. Extensive experiments demonstrate that our method achieves state-of-the-art performance across multiple datasets both qualitatively and quantitatively in zero-shot settings, as well as strong generalization to real-world indoor and outdoor scenarios.
@@ 1 Introduction
Video stereo matching is a task to reconstruct a 3D scene from a stereo camera. It estimates disparity by identifying the horizontal correspondences between rectified left and right images at each time step. Consistent and accurate disparity estimation is crucial for reconstructing and understanding real-world 3D scenes, benefiting various downstream tasks such as autonomous driving, robotic navigation, and virtual/augmented reality.
@@
To address these limitations, we propose Stereo Any Video, a robust framework for video-based stereo matching without relying on camera pose or optical flow. Inspired by recent findings in video generation [33, 48], which emphasize the role of robust feature representations in ensuring temporal coherence, we take a new perspective to design the framework for consistency: feature robustness and stability. Specifically, during feature extraction, we integrate convolutional cues with a frozen monocular video depth foundation model [7] trained on million-scale data. This ensures robust feature maps with rich priors. For correspondence matching, we introduce an all-to-all-pair correlation, which enforces smooth and reliable matching by jointly considering bidirectional correspondences from both target and reference views. For cost aggregation, we propose temporal convex upsampling within a fully temporal-based recurrent unit, playing a key role in maintaining temporal consistency. As shown in Figure 1, our method demonstrates a strong ability to produce accurate and consistent disparities, benefiting the reconstruction for dynamic scenes.
The contributions of our work can be summarized as follows: (1) Stereo Any Video, a framework that integrates monocular video depth priors to achieve accurate and temporally consistent video stereo matching. (2) A novel all-to-all-pair correlation module and a temporal convex upsampling mechanism, ensuring robust and stable cost volume representations and aggregation. (3) State-of-the-art performance on diverse benchmarks, setting a new standard for video stereo matching.
Figure 2: Overview of Stereo Any Video. Given a stereo video sequence as input, the model first extracts features and context information using trainable convolutional encoders and a frozen monocular video depth encoder (Video Depth Anything [7]). At each iteration, an all-to-all-pair correlation computes feature correlations, followed by a multilayer perception (MLP) encoder for compact representation. Disparities are iteratively refined using a 3D Gated Recurrent Unit (GRU), integrated with cost volume features, and upsampled via a temporal convex upsampling layer. This process is repeated within a cascaded pipeline to progressively recover full-resolution disparities.
@@ 2 Related Work
Video-based Stereo Matching. Video stereo matching especially in dynamic scenes presents unique challenges. Early work, such as [75], employed LSTMs with unsupervised learning for stereo videos but did not explicitly incorporate temporal information. CODD [29] introduced a per-frame stereo network, a motion network, and a fusion network designed for dynamic scenes. Neighboring-frame disparities are aligned and fused using estimated camera poses. TemporalStereo [73] extended this approach with a coarse-to-fine network that leverages past context to enhance predictions in challenging scenarios. Cheng et al. [10] developed a framework tailored for XR systems, reducing computational costs through temporal cost aggregation. TC-Stereo [70] further introduced a temporal disparity completion mechanism and a temporal state fusion module to exploit temporal information. These methods rely on auxiliary camera motion or pre-defined scene geometry to incorporate temporal information. DynamicStereo [24] introduced a transformer-based architecture that integrates temporal, stereo, and spatial attention. BiDAStereo [22] proposed an optical-flow-based alignment mechanism to propagate temporal information across the entire sequence. Building on this alignment, BiDAStabilizer [23] introduced a lightweight plug-and-play stabilizer network to improve the temporal consistency of existing image-based stereo matching methods.
@@ 3 Method
In this section, we introduce Stereo Any Video, a feed-forward video-based framework. Given a rectified stereo sequence $\{\mathbf{I}^{t}_{L},\mathbf{I}^{t}_{R}\}_{t\in(1,T)}\in\mathbb{R}^{H\times W\times 3}$ where $T$ denotes the total number of frames, $H$ and $W$ are the frame height and width, respectively, our goal is to estimate a temporally consistent sequence of disparity maps $\{\mathbf{d}^{t}\}_{t\in(1,T)}\in\mathbb{R}^{H\times W}$ aligned with the left view. As shown in Figure 2, our method consists of three key components: feature extraction (Section 3.1), correlation (Section 3.2), and temporal cost aggregation (Section 3.3).
@@ 3.1 Feature Extraction with Foundation Priors
Unlike image-based tasks, maintaining temporal consistency is the primary challenge in video stereo matching. Recent studies in video generation [33, 48] indicate that substantial variations across different frames can lead to unstable feature representations, thereby degrading temporal coherence. Therefore, ensuring feature stability across adjacent frames is crucial for achieving temporally consistent disparity estimation. Towards this, we leverage Video Depth Anything (VDA) [7] for feature extraction, which offers two key advantages: (1) trained on internet-scale data, VDA provides rich priors that can enhance disparity estimation in regions with ambiguous stereo correspondence; (2) unlike image-based monocular depth models [68, 69], VDA generates temporally stable features through an explicit temporal alignment process before the final prediction head, thereby preserving coherence across frames.
Besides, we investigate whether the depth maps from monocular video depth models can further improve stereo performance. We compare DepthCrafter [19], a representative monocular video model, with RAFTStereo [31], an image-based stereo model. Using open-source weights, we evaluate their performance on Sintel [2], as shown in Table 1 and Figure 3. Specifically, we first obtain relative depth maps from DepthCrafter (without scale and shift) and compute disparities using RAFTStereo on the same sequence. The relative depth maps are then aligned to RAFTStereo’s disparities, and objective metrics are computed against the ground truth. Additionally, we conduct a subjective user study (see supplementary material for details). As observed, while DepthCrafter produces visually plausible results—reflected in better human scores—its spatial and temporal errors are significantly larger than those of RAFTStereo. This suggests that its relative depth is “consistent but wrong”, making it unreliable for directly improving consistency within a stereo framework. Therefore, we only utilize its features as valuable priors.
@@
Table 1: Quantitative comparison of DepthCrafter [19] and RAFTStereo [31] on Sintel dataset [2].
@@
Figure 3: Qualitative comparison of DepthCrafter [19] and RAFTStereo [31] on Sintel dataset [2].
@@ 3.2 All-to-all-pair Correlation
At the $n$-th iteration for each time step, the right feature $\mathbf{F}_{R}$ is first warped toward the left feature $\mathbf{F}_{L}$ using the previous disparity ${d}_{n-1}$, producing $\mathbf{\hat{F}}_{R}$. Previous methods [31, 61] typically construct cost volumes $\mathbf{C}_{n}$ via all-pair correlation, employing a lookup mechanism to search for potential matches within a local radius. In [27], this approach was extended to an alternating local 1D and 2D search:
$\mathbf{C}_{n}(x,y)=\Big{\langle}\mathbf{F}_{L}(x,y),\hat{\mathbf{F}}_{R}(x+r_{x},y+r_{y})\Big{\rangle},$ (1)
@@
$\mathbf{C}_{n}(x,y)=\Big{\langle}\mathbf{F}_{L}(x+r_{x},y+r_{y}),\hat{\mathbf{F}}_{R}(x+r_{x},y+r_{y})\Big{\rangle}.$ (2)
@@
Table 2: Zero-shot generalization results. All methods were trained only on SceneFlow [34]. The representative image-based and video-based stereo matching models are compared on Sintel [2], Dynamic Replica [24], Infinigen SV [23], and Virtual KITTI2 [3]. The weights and parameters are fixed for evaluation. The best and second-best results are highlighted.
@@ 3.3 Temporal Cost Aggregation
At the $n$-th iteration, the all-to-all-pair correlation expands the potential matching points in $\mathbf{C}_{n}$, increasing its dimensionality from $\mathbb{R}^{H^{\prime}\times W^{\prime}\times(2r_{x}+1)(2r_{y}+1)}$ to $\mathbb{R}^{H^{\prime}\times W^{\prime}\times(2r_{x}+1)^{2}(2r_{y}+1)^{2}}$, where $H^{\prime}$ and $W^{\prime}$ denote the height and width of frame feature, respectively. We first encode $\mathbf{C}_{n}$ into a compact representation $\mathbf{E}_{n}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times L}$ using a MLP:
@@
While prior methods emphasize temporal consistency in correlation and cost aggregation, the upsampling process remains overlooked, despite its crucial role in achieving accurate and temporally consistent disparities. Towards this, we adapt the commonly used convex upsampling approach [51] to a temporal variant. In Figure 4, each high-resolution disparity pixel (yellow box) is computed as a convex combination of its $27$ low resolution neighbors across the $3$ frames (green, blue, and faint yellow boxes). The network predicts the corresponding $3\times 3\times 3$ weights using 3D convolutions, enabling a temporally consistent upsampling:
@@ 3.4 Loss Function
Recent works have introduced OPW loss [57] and TGM loss [7] in monocular video depth estimation to enhance temporal consistency. However, we found out that these losses impose a trade-off between accuracy and consistency without yielding substantial improvement in the video stereo task. Thus, we opt for a purely image-based $L_{1}$ loss while still achieving strong temporal consistency as:
@@ 4 Experiments
Table 3: Quantitative comparisons on Spring [35], Sintel [2], and KITTI Depth [54] where methods are allowed to train on any existing datasets. Model weights remain fixed during evaluation. Training datasets sharing the same domain as the test sets are highlighted. “–” represents the test set has been used in training. Abbreviations: SF – SceneFlow [34], DR – Dynamic Replica [24], ISV – Infinigen SV [23], M – Middlebury [44], KD – KITTI Depth [54], VK – Virtual KITTI2 [3]. The best and second-best results are highlighted.
    Method Training Data Scale Spring Sintel Final KITTI Depth
@@
RAFTStereo [31] 10 mixed datasets 0.6M 1.49 9.79 5.37 0.41 7.96 0.51
IGEVStereo [61] SF+M 0.04M 1.53 9.72 5.37 1.53 13.49 2.23 0.46 9.12 0.60
Selective-IGEV [55] SF+M 0.04M 1.46 9.33 5.46 1.64 13.78 2.14 0.48 9.66 0.60
DEFOM-Stereo [20] 15 mixed datasets 0.7M 1.29 9.12 4.25 0.39 7.08 0.48
MonSter [8] 10 mixed datasets 0.6M 1.25 8.65 5.19 0.41 7.73 0.60
FoundationStereo [59] 7 mixed datasets 1.3M 1.78 9.82 5.40 0.40 7.20 0.57
@@
DynamicStereo [24] SF+DR+ISV+KD 0.1M 1.07 9.64 9.36 1.38 11.84 3.38 0.42 8.21 0.57
BiDAStereo [22] SF+DR+ISV+KD 0.1M 0.90 7.90 3.92 1.33 11.34 1.97 0.42 8.33 0.58
Stereo Any Video (ours) SF+DR+ISV+VK 0.1M 0.77 6.52 3.17 0.99 9.09 1.48 0.35 6.05 0.61
@@ 4.1 Benchmark Datasets and Metric
Dataset. Our work focuses on videos captured with moving cameras, making commonly used image benchmarks unsuitable, such as Middlebury [44], ETH3D [45], and KITTI 2012/2015 [13] (two frames per scene). To evaluate our method, we use five synthetic and two real-world stereo video datasets, all featuring dynamic scenes. Sintel [2] is a synthetic dataset comprising movie sequences available in both clean and final passes. Spring [35] is a high resolution synthetic dataset. Dynamic Replica [24] is a synthetic indoor dataset featuring non-rigid objects, such as animals and people. Infinigen SV [23] is a synthetic dataset focusing on outdoor natural environments. Virtual KITTI2 [3] is a synthetic dataset simulating outdoor driving scenarios. KITTI Depth [54] is a real-world outdoor dataset captured for autonomous driving applications, with sparse depth maps obtained from a LiDAR sensor. South Kensington SV [23] is a real-world dataset capturing daily scenarios from a stereo camera for qualitative evaluation.
@@ 4.3 Evaluation
Table 4: Ablation study on the foundation priors, correlation, cost volume encoder, convex upsampling, attention, and inference settings of our model on Sintel [2] and Dynamic Replica [24]. The approach used in our final model is underlined.